OpenStudy (anonymous):

how do you solve this by using a proportion? x-2/x+3 = 3/8? help por favor/:

OpenStudy (anonymous):

\[\frac{x-2}{x+3}=\frac{3}{8}\]cross and time \[(x-2)*8=(x+3)*3\]\[8x-16=3x+9\]\[8x-3x=9+16\]\[5x=15\]\[x=\frac{15}{5}\]\[x=5\]

OpenStudy (anonymous):

ahh this helped soo much! thank you(:

OpenStudy (anonymous):

x-2/x+3=3/8 using componendo and dividendo we have x-2+x+3/x-2-x-3=3+8/3-8 which implies 2x+1/(-5)=11/(-5) or 2x+1=11 or2x=10 therefore x=5